### v1

@@ Abstract
Graph-Level Anomaly Detection (GLAD) aims to distinguish anomalous graphs within a graph dataset. However, current methods are constrained by their receptive fields, struggling to learn global features within the graphs. Moreover, most contemporary methods are based on spatial domain and lack exploration of spectral characteristics. In this paper, we propose a multi-perspective hybrid graph-level anomaly detector namely GLADformer, consisting of two key modules. Specifically, we first design a Graph Transformer module with global spectrum enhancement, which ensures balanced and resilient parameter distributions by fusing global features and spectral distribution characteristics. Furthermore, to uncover local anomalous attributes, we customize a band-pass spectral GNN message passing module that further enhances the model’s generalization capability. Through comprehensive experiments on ten real-world datasets from multiple domains, we validate the effectiveness and robustness of GLADformer. This demonstrates that GLADformer outperforms current state-of-the-art models in graph-level anomaly detection, particularly in effectively capturing global anomaly representations and spectral characteristics.
@@ 1 Introduction
In recent years, Graph Neural Networks (GNNs) have achieved significant advancements in graph representation learning, like node classification [19], link prediction [6], and graph classification [41]. Specifically, GNNs encode intricate structural and attribute information of graphs into vectors for learning within the representation space. To date, a multitude of GNN-based models have been proposed for anomaly detection in graph-structured data [10], while they predominantly focus on detecting anomalous nodes or edges [40] within a single large graph. In contrast, the domain of graph-level anomaly detection is yet to be extensively explored.
@@
Anomalous graphs manifest as outliers and may arise in various scenarios, including local attribute anomalies, substructure anomalies, global interaction anomalies [28, 32], etc. Taking biochemistry molecular data [31] as an example, Fig. 1(a) illustrates the toxic molecule of Methyl Isocyanate ($CH_{3}NCO$), where the nitrogen atom serves as a critical toxic factor. Fig. 1(b) depicts a type of muscarinic molecule, featuring a cyclic peptide structure that distinguishes it from other compounds. Fig. 1(c) showcases the main active ingredient in Nux vomica, which is Strychnine ($C_{21}H_{22}N_{2}O_{2}$), a toxic ketone alkaloid. However, as Strychnine lacks distinctive elements or specific substructures, researchers need to discern it from the complete molecular structure. Therefore, for a more comprehensive identification of various anomalous graphs, it is imperative to design models that consider both local attributes and structural information while incorporating global interactions within the graph.
Figure 1: Samples for different kinds of graph-level anomalies and differences in spectral characteristic distribution between normal and abnormal samples.
@@
The mainstream GLAD models primarily employs GNNs to jointly encode node attributes and topological structural features for acquiring node representations [49]. Additionally, they design various graph pooling functions [23] to generate graph-level representations for anomalous graphs identification. However, these methods still face several challenges. On the one hand, traditional GNN models exhibit restricted receptive fields [22], focusing solely on local neighbors or subgraph information of the current node, thereby lacking the ability to capture long-distance information interactions and global features. Although some methods attempt to learn intra-graph and inter-graph knowledge by maintaining a repository of anomalous node or graph candidates [28], they still fall short in capturing the global information within the graphs. On the other hand, when learning local features, spatial-domain GNNs tend to overlook essential high-frequency information and underlying semantic details due to their low-pass filtering characteristics [4]. Moreover, the current graph-level pooling mechanisms exhibit constrained generalization capabilities, leading to erratic performances across disparate datasets.
@@
To address the aforementioned challenges, we propose a novel graph-level anomaly detection model called GLADformer, which combines spectral-enhanced global perception and local multi-frequency information guidance. Our method, GLADformer, primarily consists of two key modules: the Spectrum-Enhanced Graph Transformer module and the Local Spectral Message Passing module. Specifically, we first introduce a Graph Transformer operator in the spatial domain, where effective graph-induced biases such as node degrees and structural information are jointly inputted, incorporating explicit spectral distribution deviations to capture anomaly information from a global perspective. Subsequently, to better explore local features and mitigate the limitations of spatial Graph Neural Networks (GNNs), we design a novel Wavelet Spectral GNN that learns discriminative local attributes and structural features in a complex spectral space. Finally, to overcome class imbalance issues and excessive confidence of traditional cross-entropy loss, we propose an improved variation-optimized cross-entropy loss. Our main contributions can be summarized as follows:
@@
    To better alleviate the issue of class imbalance and overcome the limitations of using cross-entropy as a measure for classification, we propose a weighted Variation-optimized Cross-entropy loss function.
@@
    Comprehensive experiments on a variety of datasets across ten baselines demonstrate that GALDformer exhibits competitive performance in terms of both effectiveness and efficiency.
@@ 2.1 Graph-Level Anomaly Detection
In recent years, graph anomaly detection has garnered extensive attention across various domains. However, most existing approaches focus on detecting anomalous nodes or edges within individual graph data [45, 44], while graph-level anomaly detection remains largely unexplored.
GLAD aims to differentiate deviant structures or abnormal properties within a single graph to identify anomalous graphs that exhibit substantial differences compared to the majority in a collection. State-of-the-art end-to-end methods leverage powerful GNN backbones and incorporate advanced strategies to learn graph representations suitable for anomaly detection. For instance, GLocalKD [27] and OCGTL [33] respectively combine GNN with knowledge distillation and one-class classification to detect anomalous graphs. iGAD [49] introduces an abnormal substructure-aware deep random walk kernel and node-aware kernel to capture both topological and node features. To better explore inter-graph information, GmapAD [28] maps individual graphs to a representation space by computing similarity between graphs and inter-graph candidate nodes, achieving high discriminability between abnormal and normal graphs. Some recent works focus on interpretability analysis of graph-level anomaly detection and have achieved promising results. For instance, SIGNET [25] measures the anomaly of each graph based on cross-view mutual information [12] and extracts bottleneck subgraphs in a self-supervised manner to provide explanations for anomaly discrimination. However, existing methods mostly originate from a spatial perspective, and there is still a lack of exploration regarding the influence of graph-level spectrum energy information.
@@ 2.2 Graph Transformer
Transformer [38, 42] has achieved overwhelming advantages in the NLP [7, 43] and CV [9] domains, and recently many researchers have been devoted to extending Transformer to the study of graph-structured data [48]. One of the strengths of Transformer is its ability to capture global receptive fields, but it lacks the capability to capture positional information, which poses significant limitations in graph data [37]. Recently, researchers investigate the use of Position Encoding (PE) and Structure Encoding (SE) [24, 50] within the graph domain to capture various types of graph structure features, leveraging techniques such as shortest path proximity [26] or spectral information to enhance inductive bias. For example, Graphformer [48] designs novel structural position encoding that outperforms popular GNN models in a wide range of graph prediction tasks. SAN [20] employs both sparse and global attention mechanisms at each layer and introduces learnable Position Encoding (LapPE) to replace static Laplacian eigenvectors. Exphormer [34] explores sparse attention mechanisms with virtual global nodes and extended graphs, showcasing linear complexity and desirable theoretical properties.
@@ 4 Method
Figure 2: The above image presents the overview of our model GLADformer, where (a) and (b) demonstrate the the spectrum-enhanced graph-transformer module and the local spectral message-passing module respectively.
@@ 4.1 Spectrum-enhanced Graph Transformer Module
Then the RRWP initial positional encoding for each pair of nodes is illustrated as $R_{i,j}=[I,R,R^{2},\ldots,R^{T-1}]_{i,j}\in R^{T}$. To make full use of the structural information, we utilize the position encoding $R_{ij}$ as the edge feature $e_{ij}$. Further, we cleverly combined Transformer’s self-attention mechanism and custom edge features to design a novel Graph Transformer architecture. Specifically, in the stage of calculating attention score, we replace the scalar product of Query and Key with the vector product [47], and inject the initialized edge features to obtain learnable edge features [13]. This learnable edge feature is subsequently utilized in the calculation of attention score and output encoding. The specific formula is as follows:
@@
where $\text{Norm}(\cdot)$ indicates layer-norm function, RRWP-MHA is the multi-head attention mechanism we designed before, and FFN [38] is a feed-forward network. Afterward, we merge the encoding representations from each layer of the super-node to obtain the graph-level representations.
@@
4.1.1 Spectrum Enhancement
Subsequently, as discussed in section 3.2, we have already known the positive correlation between Rayleigh entropy and high-frequency components. According to [8], the spectral energy distributions of normal and anomalous graphs are distinct in the graph spectral domain. Therefore, it is crucial to incorporate spectral energy distributions into the GLAD task. However, accurately calculating the ratios of spectral energies requires the eigenvalue-eigenvector decomposition of the graph Laplacian matrix, which has a complexity of $O(n^{3})$ and is computationally expensive for large-scale graphs. To avoid high time complexity, we employ the Rayleigh quotient $R(L,X)=\frac{X^{T}LX}{X^{T}X}$ as a substitute. For signal $X\in R^{N\times d}$, it can be proven that the diagonal elements of the Rayleigh quotient $R(L,X)$ correspond to the spectral characteristics of the signal $X$ in each feature dimension on the graph.
Given the potential sparsity of graph adjacency matrices and graph signals, as well as the possibility of high discretization levels, we utilize a single-layer MLP to derive the Rayleigh quotient vector for each graph. This vector denotes the explicit Rayleigh quotient feature and it primarily focuses on the global spectral distribution characteristics of the graph. Furthermore, we utilize this vector to enhance the aforementioned graph-level representations.
$H_{gt}^{G}=\text{MLP}\left(\text{CONCAT}\left(H^{G}_{sup},\text{MLP}\left(diag\left(R(X,L)\right)\right)\right)\right).$ (7)
@@ 4.2 Local Spectral Message-passing Module
To devise an adaptive local message passing filter, we introduce the Beta-distributed wavelet basis, which conforms to the Hammond’s graph wavelet theory [3] and is band-pass in nature. The underlying probability distribution function is $f_{Beta}\left(w\right)=w^{\alpha}(1-w)^{\beta}/B(\alpha+1,\beta+1)$, where $w\in\left[0,1\right]$, and $B(\alpha+1,\beta+1)=\alpha!\beta!/(\alpha+\beta+1)!$ is a constant. According to [36], since the eigenvalues of the normalized Laplacian matrix satisfy $\lambda\in\left[0,2\right]$, we utilize a variation like $f_{Beta}^{\ast}(\lambda)=\frac{1}{2}f_{Beta}(\frac{\lambda}{2})$. The obtained band-pass filter is expressed below:
@@
Upon meticulous analysis, it is ascertained that each constituent of the Beta band-pass wavelet kernel manifests as an amalgam of elevated powers of the adjacency matrix $A$ and the Laplacian matrix $L$. This architecture conspicuously omits specialized low-pass and high-pass filtering kernels imperative for mitigating lower and higher frequencies, potentially precipitating the omission of pivotal information. To rectify this shortfall, we propose the adoption of the generalized Laplacian low-pass and high-pass filters:
@@
where $\psi\in\left[0,1\right]$ plays a pivotal role in modulating the characteristics of both low-pass and high-pass filter kernels. Subsequently, the dual filter kernels are concurrently utilized to discern and assimilate the unadulterated low-frequency and high-frequency signals aggregating from neighbors, thereby enhancing the model’s fidelity to local structures.
@@
where READOUT function can be achieved by permutation invariant graph pooling functions like summation or mean. Finally, we fuse the features obtained from modules A and B to obtain the final graph-level representation:
@@ 4.3 Variation-optimize Cross-entropy Loss Function
The cross-entropy loss is widely used in various classification tasks and is often employed as the loss function for designing anomaly detection. However, cross-entropy suffers from phenomena such as overconfidence and struggles to adapt to scenarios like data imbalance. It is well known that the conventional evaluation metric for classification problems is accuracy, but cross-entropy is not a smooth approximation of accuracy, which significantly impacts the accuracy of model predictions. For instance, when the predicted probability of training samples is very low, cross-entropy tends to yield a tremendously large loss, even though these data points are likely to be noisy. This phenomenon is particularly pronounced in anomaly detection data, leading to overfitting of the model to noise data. To address these issues and cater to the graph anomaly detection scenario, we propose a novel cross-entropy loss function.
@@ 5 Experiment
In this section, we perform thorough experiments to validate the effectiveness of our proposed GLADformer method against ten baselines on ten real-world datasets. Furthermore, we conduct comprehensive ablation tests on GLADformer and perform visual analysis of key components.
@@ 5.1 Experiment Settings
Datasets. We conduct experiments on 10 real-world graph datasets from two popular application domains: (i) BZR, AIDS, COX2, NCI1, ENZYMES, and PROTEINS are collected from biochemistry [31], samples from minority or truly abnormal classes are considered as anomalies. Following [28], the selected anomaly samples are downsampled, retaining only 10% of the samples. (ii) MCF-7, MOLT-4, SW-620, and PC-3 are collected from PubChem [18], and they reflect the anti-cancer activity test results of a large number of compounds on cancer cell lines. Chemical compounds that exhibit antibody activity against cancer are labeled as abnormal graphs, while others are labeled as normal graphs. The statistics of these graph datasets are summarized in Table  1.
Table 1: Statistics of the 10 Datasets.
Dataset $N.G$ $N.A$ $Ratio\%$ $AVG.n$ $AVG.e$ $Attr.$
@@
Comparison Method. To illustrate the effectiveness of our proposed model, we conduct extensive experiments between GLADformer and 9 competitive methods, which can be classified into three groups: (i) Spatial GNNs with average pooling function: GCN [19], GAT [39] and GIN [46]. (ii) Graph classification methods: SAGPool [21] and GMT [2]. (iii) State-of-the-art deep GLAD methods: GLocalKD [27], iGAD [49], HimNet [32], and GmapAD [28].
@@
Experimental Settings. For our model, the dimensions of the hidden layers and output features for all three modules are set to 128 and 32 respectively. During training, the hyperparameters $T,L,M,K$ are set to 4, 6, 3 and 4 separately, and the batch size is 128 for all datasets. We use Adam as the optimizer with a learning rate of 0.001 and utilize cross-entropy loss as the basic loss function. Each dataset is randomly shuffled and split for training, validation, and testing with ratios of 70%, 15%, and 15%. For GCN, GAT and GIN, we utilize 2 layers and apply average pooling function to obtain graph-level representations. For other baselines, we use their published settings unless the parameters are specially identified in the original paper. All experiments in this paper are conducted on an NVIDIA A100-PCIE-40GB.

### v2

@@ Abstract
Graph-Level Anomaly Detection (GLAD) aims to distinguish anomalous graphs within a graph data set. However, current methods are constrained by their receptive fields, struggling to learn global features within the graphs. Moreover, most these methods are based on spatial domain and lack exploration of spectral characteristics. In this paper, we propose a multi-perspective hybrid graph-level anomaly detector named GLADformer, consisting of two key modules. Specifically, we first design a Graph Transformer module with global spectrum enhancement, which ensures balanced and resilient parameter distributions by fusing global features and spectral distribution characteristics. Furthermore, to explore local anomalous attributes, we customize a band-pass spectral GNN message passing module that enhances the model’s generalization capability. Through comprehensive experiments on ten real-world datasets from multiple domains, we validate the effectiveness and robustness of GLADformer. This demonstrates that GLADformer outperforms current state-of-the-art graph-level anomaly detection methods, particularly in effectively capturing global anomaly representations and spectral characteristics.
@@ 1 Introduction
In recent years, Graph Neural Networks (GNNs) have achieved significant advancements in graph representation learning, like node classification [17], link prediction [5], and graph classification [40], etc. Specifically, GNNs encode intricate structure and attribute information of graphs into vectors for learning within the latent representation space. To date, a multitude of GNN-based models have been proposed for anomaly detection in graph-structured data [9], while they predominantly focus on detecting anomalous nodes or edges [39] within a single large graph. In contrast, the domain of graph-level anomaly detection is yet to be extensively explored.
@@
Anomalous graphs manifest as outliers and may arise in various scenarios, including local attribute anomalies, substructure anomalies, global interaction anomalies [27, 31], etc. Taking biochemistry molecular data [30] as an example, Fig. 1 (a) illustrates the toxic molecule of Methyl Isocyanate ($CH_{3}NCO$), where the nitrogen atom serves as a critical toxic factor. Fig. 1 (b) depicts a type of muscarinic molecule, featuring a cyclic peptide structure that distinguishes it from other compounds. Fig. 1 (c) displays the main active ingredient in Nux vomica, which is Strychnine ($C_{21}H_{22}N_{2}O_{2}$), a toxic ketone alkaloid. However, as Strychnine lacks distinctive elements or specific substructures, researchers need to discern it from the complete molecular structure. Therefore, for a more comprehensive identification of various anomalous graphs, it is imperative to design models that consider both local attributes and structural information while incorporating global interactions within the graph.
Figure 1: Examples for different categories of graph-level anomalies.
@@
The mainstream GLAD models primarily employs GNNs to jointly encode node attributes and topological structural features for acquiring node representations [49]. Additionally, they design various graph pooling functions [21] to generate graph-level representations for identification of anomalous graphs. However, these methods still face several challenges. On the one hand, traditional GNNs exhibit restricted receptive fields [20], focusing solely on local neighbors or subgraph information of the current node, thereby lacking the ability to capture long-distance information interactions and global features. Although some methods attempt to learn intra-graph and inter-graph knowledge by maintaining a repository of anomalous node or graph candidates [27], they still fall short in capturing the global information within the graphs. On the other hand, when learning local features, spatial-domain GNNs tend to overlook essential high-frequency information and underlying semantic details due to their low-pass filtering characteristics [4]. Moreover, the current graph-level pooling mechanisms exhibit constrained generalization capabilities, leading to erratic performances across disparate datasets.
@@
To address the aforementioned challenges, we propose a novel graph-level anomaly detection model, which combines spectral-enhanced global perception and local multi-frequency information guidance. Our method, GLADformer, primarily consists of two key modules: the Spectrum-Enhanced Graph Transformer module and the Local Spectral Message Passing module. Specifically, we first introduce a Graph Transformer operator in the spatial domain, where effective graph-induced biases such as node degrees and structural information are jointly inputted, incorporating explicit spectral distribution deviations to capture anomaly information from a global perspective. Subsequently, to better explore local features and mitigate the limitations of spatial GNNs, we design a novel wavelet spectral GNN to learn discriminative local attributes and structural features in a complex spectral domain. Finally, to overcome class imbalance issues and excessive confidence of traditional cross-entropy loss, we propose an improved variation-optimized cross-entropy loss function. Our main contributions can be summarized as follows:
@@
    To better alleviate the issue of class imbalance and overcome the limitations of using cross-entropy as a measure for anomaly detdection, we propose a weighted variation-optimized cross-entropy loss function.
@@
    Comprehensive experiments on a variety of datasets across ten baselines demonstrate that GALDformer exhibits competitive performance in terms of both effectiveness and robustness.
@@ 2.1 Graph-Level Anomaly Detection
In recent years, graph anomaly detection has garnered extensive attention across various domains. However, most existing approaches focus on detecting anomalous nodes or edges within an individual graph [44, 43], while graph-level anomaly detection remains largely unexplored.
GLAD aims to differentiate deviant structures or abnormal properties within a single graph to identify anomalous graphs that exhibit substantial differences compared to the majority in a collection. State-of-the-art end-to-end methods leverage powerful GNN backbones and incorporate advanced strategies to learn graph representations suitable for anomaly detection. For instance, GLocalKD [26] and OCGTL [32] respectively combine GNN with knowledge distillation and one-class classification to detect anomalous graphs. iGAD [49] introduces an abnormal substructure-aware deep random walk kernel and a node-aware kernel to capture both topology information and node features. To better explore inter-graph information, GmapAD [27] maps individual graphs to a representation space by computing similarity between graphs and inter-graph candidate nodes, achieving high discriminability between abnormal and normal graphs. Some recent works focus on interpretable analysis of graph-level anomaly detection and have achieved promising results. For instance, SIGNET [23] measures the anomaly degree of each graph based on cross-view mutual information [11] and extracts bottleneck subgraphs in a self-supervised manner to provide explanations for anomaly discrimination. However, existing methods mostly originate from a spatial perspective, and there is still a lack of exploration regarding the influence of graph-level spectrum energy information.
@@ 2.2 Graph Transformer
Transformer [37, 41] has achieved overwhelming advantages in the NLP [6, 42] and CV [8, 24] domains, and recently many researchers have been devoted to extending Transformer to the study of graph-structured data [48]. One of the strengths of Transformer is its ability to capture global receptive fields, but it lacks the capability to capture positional information, which poses significant limitations in graph data [36]. Recently, researchers investigate the use of Position Encoding (PE) and Structure Encoding (SE) [22, 50] within the graph domain to capture various types of graph structure features, leveraging techniques such as shortest path proximity [25] or spectral information to enhance inductive bias. For example, Graphformer [48] designs novel structural position encoding that outperforms popular GNN models in a wide range of graph prediction tasks. Further, SAN [18] employs both sparse and global attention mechanisms at each layer and introduces learnable Position Encoding (LapPE) to replace static Laplacian eigenvectors. Exphormer [33] explores sparse attention mechanisms with virtual global nodes and extended graphs, showcasing linear complexity and desirable theoretical properties.
@@ 4 Method
Figure 2: The above image presents the overview of our model GLADformer, where (a) and (b) demonstrate the the spectrum-enhanced graph-transformer module and the local spectral message passing module respectively.
@@ 4.1 Spectrum-enhanced Graph Transformer Module
Then the RRWP initial positional encoding for each pair of nodes is illustrated as $R_{i,j}=[I,R,R^{2},\ldots,R^{T-1}]_{i,j}\in R^{T}$. To make full use of the structural information, we utilize the position encoding $R_{ij}$ as the edge feature $e_{ij}$. Further, we cleverly combine Transformer’s self-attention mechanism and customized edge features to design a novel Graph Transformer architecture. Specifically, in the stage of calculating attention score, we replace the scalar product of Query and Key with the vector product [46], and inject the initialized edge features to obtain learnable edge features. This learnable edge feature is subsequently utilized in the calculation of attention score and output encoding. The specific formula is as follows:
@@
where $\text{Norm}(\cdot)$ indicates layer-norm function, RRWP-MHA is the multi-head attention mechanism designed before, and FFN [37] is a feed-forward network. Afterward, we merge the encoding representations from each layer of the super-node to obtain the graph-level representations.
@@
4.1.1 Spectrum Enhancement.
Subsequently, as discussed in section 3.2, we have already known the positive correlation between Rayleigh quotient and high-frequency components. According to [7], the spectral energy distributions of normal and anomalous graphs are distinct in the graph spectral domain. Therefore, it is crucial to incorporate spectral energy distributions into the GLAD task. However, accurately calculating the ratios of spectral energies requires the eigenvalue-eigenvector decomposition of the graph Laplacian matrix, which has a complexity of $O(n^{3})$ and is computationally expensive for large-scale graphs. To avoid high time complexity, we employ the Rayleigh quotient $R(L,X)=\frac{X^{T}LX}{X^{T}X}$ as a substitute. For signal $X\in R^{N\times d}$, it can be proven that the diagonal elements of the Rayleigh quotient $R(L,X)$ correspond to the spectral characteristics of the signal $X$ in each feature dimension on the graph.
Given the potential sparsity of graph adjacency matrices and graph signals, as well as the possibility of high discretization levels, we utilize a single-layer MLP to derive the Rayleigh quotient vector for each graph. This vector denotes the explicit Rayleigh quotient feature and it primarily focuses on the global spectral distribution characteristics of the graph. Then, we utilize this to enhance the previously obtained graph-level representation.
$H_{rq}^{G}$ $=\text{MLP}\left(diag\left(R(X,L)\right)\right),$ (7)
$H_{gt}^{G}=$ $\text{MLP}\left(\text{CONCAT}\left(H^{G}_{sup},H_{rq}^{G}\right)\right).$
@@ 4.2 Local Spectral Message-passing Module
To devise an adaptive local message passing filter, we introduce the Beta-distributed wavelet basis [35], which conforms to the Hammond’s graph wavelet theory [3] and is band-pass in nature. The underlying probability distribution function is $f_{Beta}\left(w\right)=w^{\alpha}(1-w)^{\beta}/B(\alpha+1,\beta+1)$, where $w\in\left[0,1\right]$, and $B(\alpha+1,\beta+1)=\alpha!\beta!/(\alpha+\beta+1)!$ is a constant. According to [35], since the eigenvalues of the normalized Laplacian matrix satisfy $\lambda\in\left[0,2\right]$, we utilize a variation like $f_{Beta}^{\ast}(\lambda)=\frac{1}{2}f_{Beta}(\frac{\lambda}{2})$. The obtained band-pass filter is expressed below:
@@
Upon meticulous analysis, it is ascertained that each constituent of the Beta band-pass wavelet kernel manifests as an amalgam of elevated powers of the adjacency matrix $A$ and the Laplacian matrix $L$. This architecture conspicuously lacks specialized low-pass and high-pass filtering kernels imperative for mitigating lower and higher frequencies, potentially precipitating the omission of pivotal information. To rectify this shortfall, we propose the adoption of the generalized Laplacian low-pass and high-pass filters:
@@
where $\psi\in\left[0,1\right]$ plays a pivotal role in modulating the characteristics of both low-pass and high-pass filter kernels. Subsequently, the dual filter kernels are concurrently utilized to discern and assimilate the pure low-frequency and high-frequency signals aggregating from neighbors, thereby enhancing the model’s fidelity to local structures.
@@
where READOUT function can be achieved by permutation invariant graph pooling functions like summation or mean [27]. Finally, we fuse the features obtained from two modules to obtain the final graph-level representation:
@@ 4.3 Variation-optimize Cross-entropy Loss Function
The cross-entropy loss is widely used in various classification tasks and is often employed as the loss function for designing anomaly detection. However, cross-entropy suffers from phenomena such as overconfidence and struggles to adapt to scenarios like data imbalance. It is well known that the conventional evaluation metric for classification problems is accuracy, but cross-entropy is not a smooth approximation of accuracy, which significantly impacts the precision of model predictions. For instance, when the predicted probability of training samples is very low, cross-entropy tends to yield a tremendously large loss, even though these data points are likely to be noises. This phenomenon is particularly pronounced in anomaly detection data, leading to overfitting of the model to noise data. To address these issues and cater to the graph anomaly detection scenarios, we propose a novel cross-entropy loss function.
@@ 5 Experiment
In this section, we perform thorough experiments to validate the effectiveness of our proposed GLADformer method against nine baselines on ten real-world datasets. Furthermore, we conduct comprehensive ablation tests on GLADformer and perform visual analysis of key components.
@@ 5.1 Experiment Settings
Datasets. We conduct experiments on 10 real-world graph datasets from two popular application domains: (i) BZR, AIDS, COX2, NCI1, ENZYMES, and PROTEINS are collected from biochemistry [30], samples from minority or truly abnormal classes are considered as anomalies. Following [27], the selected anomaly samples are downsampled, retaining only 10% of the samples. (ii) MCF-7, MOLT-4, SW-620, and PC-3 are collected from PubChem [16], and they reflect the anti-cancer activity test results of a large number of compounds on cancer cell lines. Chemical compounds that exhibit antibody activity against cancer are labeled as abnormal graphs, while others are labeled as normal graphs. The statistics of these graph datasets are summarized in Table 1.
Table 1: The statistics of the 10 datasets.
Dataset $N.G$ $N.A$ $Ratio\%$ $AVG.n$ $AVG.e$ $Attr$
@@
Comparison Method. To illustrate the effectiveness of our proposed model, we conduct extensive experiments between GLADformer and 9 competitive methods, which can be classified into three groups. (i) Spatial GNNs with average pooling function: GCN [17], GAT [38] and GIN [45]. (ii) Graph classification methods: SAGPool [19] and GMT [2]. (iii) State-of-the-art deep GLAD methods: GLocalKD [26], iGAD [49], HimNet [31], and GmapAD [27].
@@
Experimental Settings. For our model, the dimensions of the hidden layers and output features for all three modules are set to 128 and 32 respectively. During training, the hyperparameters $T,L,M,K$ are set to 4, 6, 3 and 4 separately, and the batch size is 128 for all datasets. We use Adam as the optimizer with a learning rate of 0.001 and utilize cross-entropy loss as the basic loss function. Each dataset is randomly shuffled and split for training, validation, and testing with ratios of 70%, 15%, and 15%. For GCN, GAT and GIN, we utilize 2 layers and apply average pooling function to obtain graph-level representations. For other baselines, we use their published settings unless the parameters are specially identified in the original paper. All experiments in this work are conducted on an NVIDIA A100-PCIE-40GB.